Home‎ > ‎Reports‎ > ‎Crew101‎ > ‎

day13


Crew 101 Log Book for March 24, 2011

Photos of the day





Commander's Check-In Report

Tom Caillouet Reporting

Crew Health: Quite good

EVAs and Times: None

Number of Reports: 5 (Commander, Engineering, Journalist, Chef, Science)

Field Results: N/A

Plans for Tomorrow: Breaking sim to retrieve the repeater.


Commander's Report

Tom Caillouet Reporting

It has been another day of science for Crew 101. Most of the crew spent the day in the lab prepping samples and completing various analyses (I'm not exactly sure what it all entails, I am but a mere engineer). Cody and I spent some time attempting to get the ATVs to cooperate, but when our repair methodology is attempting to jump the non-working units, we run out of strategies fairly quickly. The remainder of the day has been spent wrapping up our work here. Writing reports (though full science results won't be ready until tomorrow), cleaning, if we can't gather any more samples, we'll at least have am extremely smooth transition ready for the next crew's arrival. Tomorrow, we will officially break sim to bring down the radio repeater - no need to endanger more equipment.


Engineering Report

Cody Hall Reporting

Generator/Electricity:

• Kitty (Diesel Generator) (Used? Yes/No): Yes

• Honey (Gas Generator) (Used? Yes/No): No

• Battery Charge Level:

o V everstart

o V armor plate

o V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run

Fuel Status:

• Diesel (Full/0.75/0.5/0.25/Empty): 0.75

• Propane Tank (Full/0.75/0.5/0.25/Empty): 0.25

• Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.5

Notes/Comments:

Oil:

• Kitty Oil Quantity (# of Quarts in storage): Changed/replaced by DG

• Honey Oil Quantity (# of Quarts in storage): Leave it to DG

• ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

• Car Oil Quantity (# of Quarts in storage): Leave it to DG

Notes/Comments:

Water Status:

• Outside Potable Water Tank Level (inches from bottom): 13

• Trailer Potable Water Tank Level (inches from bottom): full

• Hab Potable Water Tank Level (inches from bottom): 5

• Potable Water Meter Reading: 438436

Notes/Comments :


GreenHab:

• Condition of plants in Tank 1: unchanged

• Condition of plants in Tank 2: unchanged

• Condition of plants in Tank 3: water level dropped another ½ inch

• Greenhab notes/ comments: straining tank is draining, but is still not down to sand. Only dropped another inch

Observatory:

• Telescope Used? (Yes/no) No

Observatory Notes: DG stopped by to install new hand control

Transportation:

Hab Car (NOT RENTAL)

• Used (yes/no): yes

• Oil Checked (yes/no): No

• Tire Status: Not checked, good

• Odometer at end of day: 123701.7

• Notes/Comments on Hab Car:

ATV 1 (Opportunity):

• Used: yes

• Oil Checked: no

• Fuel Consumed: replaced

• Tire Status: ok

• Comments:


ATV 2 (Spirit):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:


ATV 3 (Viking I):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments: DG came by and started it. We apparently were not pulling the cord hard enough.


ATV 4 (Viking II):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments: DG thinks it has a real problem, and he’ll look at it later. For now, this one doesn’t work.

ATV Comments:

Heating and Ventilation:

• Thermostat setting upstairs (Farenheit): 68

• Actual temperature upstairs (Farenheit): 70

• Thermostat setting downstairs (Farenheit): 67

• Actual temperature downstairs (Farenheit): 65

Notes/Comments:


Computers/Networking Infrastructure: working

Notes/comments on internet/computers:


General Engineering:

The porthole to the outer door of the EVA airlock was found outside today. We assume it came out of the door because of the wind. It is now sitting in the EVA prep room. DG was told about this.


Chef's Report

Christine Redmond Reporting

1a) Today was a pre-prepared.

1b) Was today a special day celebrated at the main meal with special

food or activity?

No

2) List the foods served at today's main meal (usually the evening

meal), giving full names of each.

Chili

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

None

4) Recipes for anything more complicated than following package

directions exactly, or rehydrating.

5) Lessons learned (if any)

6) Comments/questions for Kim and Jean, the food study investigators


Journalist Report

Christine Redmond Reporting

Happy Thursday earthlings! Crew 101 has only two more days on Mars. Although we are all going to miss the red planet, we are all excited for burgers in our near future. Last night Crew 101 enjoyed a nice clear night to stargaze. That was definitely a night sky you can only see on Mars! We will all hold on to that memory for a long time.

Crew 101 had a busy and rewarding day starting to wrap up our work here on Mars. Christina, Mitesh and I had a productive morning in the lab. I ran some tests for our contamination study and Mitesh and Christina were busy running some biology tests. Cody and Tommy worked this afternoon to try to get the ATVs back up and running. Unfortunately they were not able to get them running. Today is our last day on the food study. Tonight we celebrate our last “pre-prepared day” with chili!